# HERBench: A Benchmark for Multi-Evidence Integration in Video Question Answering

### v1

@@ 7.8 Dataset Statistics
HERBench comprises 26,806 questions derived from 336 unique videos. The videos feature substantial duration (avg. 395s, range 60-2100s) to ensure temporal dispersion of evidence. Sources include HD-EPIC, WildTrack, PersonPath22, and movie trailers.
@@ 8.1 Extended MRFS Analysis
Table 4: Per-task MRFS statistics Computed with $x=16$ using Qwen2.5-VL and AKS.
@@
Figure 9: Impact of Evidential Requirement on Model Accuracy. We plot the Mean Minimum Required Frame-Set (MRFS) against Full-context Accuracy ($k=16$), measured using Qwen 2.5 VL 7B, across four video QA benchmarks. The data suggests an inverse trend: as the necessity to aggregate distinct visual cues increases (higher MRFS), model performance tends to decrease. HERBench (green) imposes a higher evidential burden ($\text{MRFS}\ 5.49$), highlighting the potential challenges current Video-LLMs face in multi-evidence integration relative to benchmarks with lower requirements like NeXT-QA.
As illustrated in Figure 9, there appears to be an inverse relationship between the evidential demand of a benchmark—quantified by the Mean MRFS—and the performance of state-of-the-art Video-LLMs. Existing benchmarks such as NeXT-QA exhibit a lower evidential requirement (2.61MRFS2.61\ $\text${MRFS}), where Qwen 2.5 VL 7B achieves relatively high accuracy ($76.3\%$), possibly due to the feasibility of single-frame shortcuts or language priors. In contrast, HERBench presents a higher burden (5.49MRFS5.49\ $\text${MRFS}), designed to require the integration of non-redundant, temporally separated cues. This increased demand coincides with a lower accuracy of $35.9\%$, a pattern that is consistent with the hypothesized fusion deficit in current architectures. These results suggest that while models may be effective at retrieving isolated frames, their capacity for compositional reasoning appears to be increasingly challenged as the number of required evidence pieces grows.
Table 5: Frame Selection Ablation. Accuracy (%) on a random subsample of questions. GT Frames (OF) represents the upper bound with manually curated evidence.
@@ 8.2 Full Frame-Selection Ablation
To more precisely disentangle the role of evidence retrieval from that of multi-evidence fusion, we perform an extensive ablation over five frame selection strategies—Uniform, Vanilla-BLIP, BOLT-ITS, AKS, and Oracle Frames (OF)—and evaluate their effect across all twelve HERBench tasks (Table 5). Overall, learned strategies such as BOLT-ITS and AKS provide moderate gains over Uniform sampling, reflecting their ability to prioritize query-relevant frames while maintaining broader temporal coverage. However, their improvements are uneven across tasks: both methods show the largest benefits in sparse-evidence settings such as [TSO] and [FAM], where the critical evidence may appear only briefly within long videos. The oracle-based setting establishes an upper bound by supplying the manually curated evidence frames used during dataset construction. As shown in the rightmost column of Table 5, all three representative models experience non-trivial but still limited performance improvements in the OF regime (typically +3+3-66 absolute accuracy points relative to the best learned selector).
@@ 9 Illustrative Examples for All Tasks
Figure 10 presents an example of the Temporal Shot Ordering (TSO) task, which requires reconstructing the chronological order of four non-overlapping shots. Figure 11 shows the Multi-Person Duration Reasoning (MPDR) task, where models must compare visible-time intervals across multiple individuals. Figure 12 illustrates the Action Sequence Integrity & Identification (ASII) task, requiring identification of the correct sequence among plausible permutations of narrated events.
@@
Figure 13 shows the Appearance-Grounded Behavior Interactions (AGBI) task, where models must track a target described only by appearance and determine who interacts with them. Figure 14 provides an example of the Appearance-Grounded Attribute Recognition (AGAR) task, requiring attribute extraction anchored to the tracked target. Figure 15 illustrates the Appearance-Grounded Localization Trajectory (AGLT) task, where the model must infer how the target enters or exits the scene.
@@
Figure 16 presents the False Action Memory (FAM) task, requiring verification of which plausible action did not occur in the video. Figure 17 shows the Scene Verification & Arrangement (SVA) task, combining faithful and perturbed shot descriptions to assess fine-grained scene-level verification and ordering. Figure 18 depicts the False Object Memory (FOM) task, requiring identification of a plausible but absent object interaction.
@@
Figure 19 provides an example of the Multi-Entities Grounding & Localization (MEGL) task, where models must verify which appearance-described individuals actually appear in the video. Figure 20 illustrates the Action Counting (AC) task, requiring enumeration of all instances of a specified action–object pair across the entire video. Finally, Figure 21 shows the Region-Localized People Counting (RLPC) task, where the model must count unique individuals entering through specific spatial regions.

### v2

@@ 7.9 Dataset Statistics
HERBench comprises 26,806 questions derived from 336 unique videos. The videos feature substantial duration (avg. 395s, range 60-2100s) to ensure temporal dispersion of evidence. Sources include HD-EPIC [35], WildTrack [5], PersonPath22 [40], and movie trailers.
@@ 8.1 Extended MRFS Analysis
Table 6: Per-task MRFS statistics Computed with $x=16$ using Qwen2.5-VL [3] and AKS [43].
@@
Figure 9: Impact of Evidential Requirement on Model Accuracy. We plot the Mean Minimum Required Frame-Set (MRFS) against Full-context Accuracy ($k=16$), measured using Qwen 2.5 VL 7B [3], across ten video QA benchmarks. The dashed line indicates the fitted trend. A clear inverse correlation emerges: as the evidential burden increases (higher MRFS), model performance tends to decrease ($R^{2}=0.82$). HERBench (green star) occupies the high-demand end of the spectrum ($\text{MRFS}\ 5.49$), highlighting the challenges current Video-LLMs face in multi-evidence integration relative to lower-requirement benchmarks such as TemporalBench and NeXT-QA.
As illustrated in Figure 9, there is a pronounced inverse relationship between the evidential demand of a benchmark—quantified by the Mean MRFS—and the performance of state-of-the-art Video-LLMs. Benchmarks with low evidential requirements, such as TemporalBench (2.21MRFS2.21\ $\text${MRFS}, $87.5\%$) and NeXT-QA [50] (2.61MRFS2.61\ $\text${MRFS}, $76.3\%$), allow Qwen 2.5 VL 7B [3] to achieve relatively high accuracy, possibly due to the feasibility of single-frame shortcuts or language priors. Mid-range benchmarks—CVRR-ES, MVBench, LongVideoBench, and MMBench-Video—cluster between 2.82.84.4MRFS4.4\ $\text${MRFS} with accuracies that progressively decline from $70.0\%$ to $50.8\%$, tracing the fitted trend closely. At the high-demand end, MINERVA (5.14MRFS5.14\ $\text${MRFS}, $37.9\%$), Video-MME (5.31MRFS5.31\ $\text${MRFS}, $50.6\%$), and HERBench (5.49MRFS5.49\ $\text${MRFS}, $35.9\%$) impose substantially greater evidential burdens, coinciding with markedly lower accuracies. Notably, AGQA (3.42MRFS3.42\ $\text${MRFS}, $44.0\%$) falls below the trend line, suggesting that factors beyond evidential density—such as compositional question complexity—can independently depress performance. The broadened comparison across ten benchmarks strengthens the evidence for a fusion deficit in current architectures: while models may be effective at retrieving isolated frames, their capacity for compositional reasoning degrades consistently as the number of required evidence pieces grows. HERBench, positioned at the extreme of this spectrum, is specifically designed to stress-test this bottleneck by requiring the integration of non-redundant, temporally dispersed cues.
Table 7: Frame Selection Ablation. Accuracy (%) on a random subsample of questions using InternVL3.5 [45], Qwen3-VL [2], and Ovis-2.5 [27] with Uniform, Vanilla-BLIP [21], BOLT-ITS [25], AKS [43], and GT Frames (OF) selectors. GT Frames represents the upper bound with manually curated evidence.
@@ 8.2 Full Frame-Selection Ablation
To more precisely disentangle the role of evidence retrieval from that of multi-evidence fusion, we perform an extensive ablation over five frame selection strategies—Uniform, Vanilla-BLIP [21], BOLT-ITS [25], AKS [43], and Oracle Frames (OF)—and evaluate their effect across all twelve HERBench tasks (Table 7). Overall, learned strategies such as BOLT-ITS and AKS provide moderate gains over Uniform sampling, reflecting their ability to prioritize query-relevant frames while maintaining broader temporal coverage. However, their improvements are uneven across tasks: both methods show the largest benefits in sparse-evidence settings such as [TSO] and [FAM], where the critical evidence may appear only briefly within long videos. The oracle-based setting establishes an upper bound by supplying the manually curated evidence frames used during dataset construction. As shown in the rightmost column of Table 7, all three representative models experience non-trivial but still limited performance improvements in the OF regime (typically +3+3-66 absolute accuracy points relative to the best learned selector).
@@
8.3 MRFS robustness across backbones and frame selectors
Table 8: MRFS robustness across models and keyframe selectors. Benchmarks: NExT-QA [50], MVBench [22], LongVideoBench [48], and HERBench.
Benchmark Model on AKS (MRFS / Acc.) Selector on Qwen2.5-VL (MRFS / Acc.)
Qwen2.5-VL Gemini2.5F LLaVA-OV-1.5 GPT-4o AKS BOLT $T^{*}$
NExT-QA 2.61 / 65.79 3.81 / 76.29 3.45 / 68.77 3.85 / 70.81 2.61 / 65.79 3.64 / 75.33 2.78 / 64.56
MVBench 3.52 / 56.71 3.92 / 57.41 3.69 / 54.24 4.09 / 54.42 3.52 / 56.71 3.73 / 56.28 3.53 / 55.53
LongVideoBench 4.07 / 41.38 4.57 / 64.59 4.50 / 61.49 4.92 / 59.01 4.07 / 41.38 4.89 / 49.14 4.11 / 48.59
[…]
To assess whether MRFS-based benchmark comparison is sensitive to the choice of backbone or selector, we evaluate MRFS under a range of configurations: four backbone models (Qwen2.5-VL [3], Gemini 2.5 Flash [7], LLaVA-OV-1.5 [1], and GPT-4o [33]) with AKS [43], and three selectors (AKS, BOLT [25], and $T^{\star}$ [54]) with Qwen2.5-VL (see Tab. 8). Here, $T^{\star}$ serves as a non-CLIP baseline. All results are computed on 50% stratified random samples from each benchmark. Across all tested settings, the benchmark ordering remains stable, with HERBench consistently yielding the highest MRFS, followed by LongVideoBench, MVBench, and NExT-QA. These results support MRFS as a robust measure of dataset-level evidential requirement.
@@ 9 Illustrative Examples for All Tasks
Figure 10 presents an example of the Temporal Shot Ordering ([TSO]) task, which requires reconstructing the chronological order of four non-overlapping shots. Figure 11 shows the Multi-Person Duration Reasoning ([MPDR]) task, where models must compare visible-time intervals across multiple individuals. Figure 12 illustrates the Action Sequence Integrity & Identification ([ASII]) task, requiring identification of the correct sequence among plausible permutations of narrated events.
@@
Figure 13 shows the Appearance-Grounded Behavior Interactions ([AGBI]) task, where models must track a target described only by appearance and determine who interacts with them. Figure 14 provides an example of the Appearance-Grounded Attribute Recognition ([AGAR]) task, requiring attribute extraction anchored to the tracked target. Figure 15 illustrates the Appearance-Grounded Localization Trajectory ([AGLT]) task, where the model must infer how the target enters or exits the scene.
@@
Figure 16 presents the False Action Memory ([FAM]) task, requiring verification of which plausible action did not occur in the video. Figure 17 shows the Scene Verification & Arrangement ([SVA]) task, combining faithful and perturbed shot descriptions to assess fine-grained scene-level verification and ordering. Figure 18 depicts the False Object Memory ([FOM]) task, requiring identification of a plausible but absent object interaction.
@@
Figure 19 provides an example of the Multi-Entities Grounding & Localization ([MEGL]) task, where models must verify which appearance-described individuals actually appear in the video. Figure 20 illustrates the Action Counting ([AC]) task, requiring enumeration of all instances of a specified action–object pair across the entire video. Finally, Figure 21 shows the Region-Localized People Counting ([RLPC]) task, where the model must count unique individuals entering through specific spatial regions.